OpenStudy (anonymous):

How can you isolate y in the equation: (D^2 + (y + d/2)^2)^1/2 = (D^2 + (y - d/2)^2)^1/2 + 10k

OpenStudy (anonymous):

\[\sqrt{(D^2 + (y + d/2)^2)} = \sqrt{(D^2 + (y - d/2)^2)} + 10k\] \[d \neq D\]